A block slides from rest from the top of an inclined plane 8 m long which is inclined 35° with  the horizontal. If the coefficient of kinetic friction is 0.20, determine how long it will take the block to reach the bottom of the plane.


1
Expert's answer
2021-03-03T11:24:18-0500

Let mass of block = m kg

Taking acceleration due to gravity = 9.8 m/s2


Coefficient of friction, "\mu=0.20"

Length of inclined plane = 8 m

From the above figure, balancing the forces perpendicular to the inclined surface,

"N=mgcos\theta\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space(1)"

Balancing the forces along the inclined plane

"ma=mgsin\theta -\mu N\\\Rightarrow ma=mgsin\theta-\mu mgcos\theta\space\space\space\space\space\space\space\space\space\space\space\space\space from(1)\\\Rightarrow a=g(sin\theta-\mu cos\theta)"


From Newton's laws of motion,

"s=ut+\dfrac{1}{2}at^2\\\Rightarrow t=\sqrt{\dfrac{2s}{a}}\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space Since\space u=0\space m/s^2"

"\Rightarrow t=\sqrt{\dfrac{2\times8}{9.8(sin35\degree-0.20cos35\degree)}}\\\Rightarrow t=1.996\space s\approx2\space s"